OpenStudy (anonymous):

PR has an endpoint at (25, -5) and a midpoint of (18, -1). What is the value of the x-coordinate of the other endpoint?

OpenStudy (anonymous):

@LaloHao can you help me

OpenStudy (anonymous):

someone or anybody

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

wait a bit pls

OpenStudy (anonymous):

okay THANKS

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ a}}\quad ,&{\color{blue}{ b}})\quad &({\color{red}{ c}}\quad ,&{\color{blue}{ d}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\\ \quad \\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 25}}\quad ,&{\color{blue}{ 5}})\quad &({\color{red}{ a}}\quad ,&{\color{blue}{ b}}) \end{array}\qquad \left(\cfrac{{\color{red}{ a}} + {\color{red}{ 25}}}{2}\quad ,\quad \cfrac{{\color{blue}{ b}} + {\color{blue}{ 5}}}{2} \right)\implies (18\ ,\ -1)\\ \quad \\ \cfrac{{\color{red}{ a}} + {\color{red}{ 25}}}{2}=18\qquad \qquad \cfrac{{\color{blue}{ b}} + {\color{blue}{ 5}}}{2}=-1\)

OpenStudy (anonymous):

Beat me to it was doing in my notebook lol

OpenStudy (anonymous):

lol THANK YOU SO MUCH